# A Stitch in Time: Learning Procedural Workflow via Self-Supervised Plackett–Luce Ranking

### v1

@@ 3.5 Image Branch: Local and Spatio-temporal Learning
Masked Image Modeling. We adopt the masked image modeling (MIM) objective from iBOT [Zhou2022IBOT:Tokenizer] on the current frame $v_{t}$ to establish a robust frame-level feature representation, optimized with the loss $\mathcal{L}_{\text{MIM}}$.
@@ 4 Experiments
1 Pairwise [Hu2021ContrastLearning] $75.8$ $75.4$
2 Permutation (CE) [Misra2016ShuffleVerification] $74.5$ $70.1$
@@ 4.1 Experimental Setup
Datasets and Evaluation Protocols. We evaluate on five widely used datasets from the surgical and cooking domains, selected for their long-range, multi-step procedural workflows with fixed sequential dependencies. For surgical procedures, we evaluate the temporal phase recognition task using AutoLaparo [Wang2022AutoLaparo:Hysterectomy] (laparoscopic hysterectomy), along with Cholec80 [Twinanda2017EndoNet:Videos] and M2CAI16 [Stauder2016TheChallenge] (both laparoscopic cholecystectomy). For cooking activities, we evaluate the temporal action segmentation task using GTEA [Fathi2011LearningActivities] (egocentric cooking) and Breakfast [Kuehne2014TheActivities] (third-person cooking). We assess feature quality using two standard protocols: Linear Probing, which trains a linear classifier on frozen features, and k-NN Classification [Caron2021EmergingTransformers], which uses $k=20$ nearest neighbors in the feature space for non-parametric evaluation. For k-NN, we report top-1 accuracy. For Linear probing, we follow established benchmarks to report frame-wise accuracy and F1-score for surgical datasets [Che2025LEMON:Settings, Wang2022AutoLaparo:Hysterectomy, Liu2025LoViT:Recognition], and frame-wise accuracy, segmental edit distance, and the segmental F1 score at overlapping thresholds 10%, 25% and 50% for cooking datasets [Li2023MS-TCN++:Segmentation, Fathi2011LearningActivities, Kuehne2014TheActivities].
@@
Pretraining Details. We adopt ViT-B/16 [Dosovitskiy2020AnScale] as our framework’s backbone. For our temporal ranking objective ($\mathcal{L}_{\text{vid}}$), we sample k=8 frames from each video of length $T$. We first select a random starting frame $t_{0}$ and then a random, uniform step length $\Delta t$, such that $t_{0}+7\Delta t\leq T$. Videos are resampled multiple times per epoch, providing broad coverage of the procedural structure. For the $\mathcal{L}_{\text{jigsaw}}$ objective, the past and future context frames are sampled from the temporal ranges [-2.5, -1.5]s and [+1.5, +2.5]s relative to the current frame, following [Liu2025WhenLearning]. For the $\mathcal{L}_{\text{MIM}}$ objective, we apply block-wise masking with a ratio of 30% as in [Zhou2022IBOT:Tokenizer]. The loss weights (Eq. 5) are set to $\lambda_{1}=1$, $\lambda_{2}=1$, and $\lambda_{3}=0.4$ with analysis in supplementary material.
Optimization Strategies. We employ different pretraining optimization strategies based on the target domain. For surgical datasets, models are pretrained for 30 epochs on the large-scale surgical video dataset LEMON [Che2025LEMON:Settings] to leverage its domain-specific knowledge. We use AdamW with a base learning rate of $4\times 10^{-4}$, a 3-epoch cosine warmup, and a total batch size of 240. For cooking datasets, we pretrain directly on their respective official training sets for 100 epochs with 10 warmup epochs, as no comparable large-scale pretraining dataset exists for this highly variable domain; other optimizer settings are kept identical. Experiments utilize 4 NVIDIA A100 (40GB) GPUs and PyTorch 2.5.1 [Paszke2019PyTorch:Library].
Figure 7: Attention maps queried by the [CLS] token comparing our method, PL-Stitch, against DINO, iBOT, and VideoMAEv2 on a sequence of frames from AutoLaparo [Wang2022AutoLaparo:Hysterectomy] during the Dividing Ligament and Peritoneum phase. The light blue box indicates the instrument operating area, which is critical for understanding the surgical phase. While other methods produce diffuse attention that changes erratically over time, our PL-Stitch maintains a consistent and accurate focus on the target area (the instrument and its operating area).
@@ 4.2 Main Quantitative Results
Surgical Phase Recognition. As shown in Table 1, PL-Stitch consistently outperforms all baselines on AutoLaparo, Cholec80, and M2CAI16 under both linear probing and k-NN evaluation. This demonstrates superior feature quality over both generalist (G) and specialist (S) methods. The performance gains are particularly striking in the k-NN evaluation, which directly assesses feature space quality. On Cholec80, PL-Stitch achieves 81.7% k-NN accuracy, a significant +11.4 pp gain over the strong iBOT baseline [Zhou2022IBOT:Tokenizer]. This trend holds for AutoLaparo (+7.2 pp) and M2CAI16 (+9.1 pp) as well. Furthermore, PL-Stitch also achieves the highest scores in all linear probing metrics, underscoring the discriminative power of its frozen features.
@@ 4.3 Ablation Study
We conduct ablations on the Cholec80 phase recognition task by pretraining on LEMON [Che2025LEMON:Settings] and report top-1 accuracy under linear probing and k-NN (k = 20). We use a ViT-S/16 backbone to reduce computational cost. Our default settings in the Tables 34, and 5 are highlighted in gray.
Effect of PL-Stitch Components. We analyze the contribution of each objective in Table 3. Our baseline (row 1), with only the $\mathcal{L}_{\text{MIM}}$ objective [Zhou2022IBOT:Tokenizer], achieves 69.4% k-NN accuracy. Adding our core temporal ranking objective ($\mathcal{L}_{\text{vid}}$) (row 2) provides the largest single boost, improving k-NN accuracy by 9.5 pp to 78.9%. Adding only the spatio-temporal jigsaw objective ($\mathcal{L}_{\text{jigsaw}}$) (row 3) provides a smaller benefit. Finally, our full PL-Stitch model (row 4), combining all three objectives, achieves the highest accuracy at 80.2%. This demonstrates that our global temporal and local jigsaw objectives are complementary, and their joint optimization leads to the most robust representation.
Impact of PL Ranking Formulation. We compare our Plackett-Luce (PL) formulation against traditional ordering objectives (Table 4). We replace our $\mathcal{L}_{\text{vid}}$ with two common baselines: 1) a Pairwise Loss [Hu2021ContrastLearning], which enforces the order of all pairs in the sequence, and 2) a Permutation Classification loss [Misra2016ShuffleVerification], which treats each shuffled permutation of the sequence as a distinct class. Both baselines perform worse than our PL objective, confirming that the probabilistic, listwise formulation of the PL model is key to learning a robust, progress-aware representation.
@@ 4.4 Qualitative Results
Feature Space Visualization. Fig. 6 visualizes the frozen Cholec80 features using t-SNE [VanDerMaaten2014AcceleratingAlgorithms]. Baseline features from VideoMAEv2, DINO, and iBOT overlap heavily, showing no clear phase separation. In contrast, PL-Stitch forms distinct, well-separated clusters that align with ground-truth phases. Its clustering quality ($\text{ARI}:0.3536,\text{NMI}:0.4546$) is more than twice as high as the second-best baseline, DINO ($\text{ARI}:0.1043,\text{NMI}:0.2051$). This clear separation visually explains its high k-NN accuracy and confirms its learned procedural awareness.
@@ A.1 Plackett-Luce Ranking Formulation
Our framework leverages the Plackett-Luce (PL) model [Plackett1975ThePermutations, Luce1977TheYears] to structure both the global and local objectives as listwise ranking problems. The PL model is parameterized by a vector of positive real-valued scores $s=(s_{1},\dots,s_{K})$, where $K$ is the number of items being ranked. The score $s_{i}$ represents the overall utility or preference strength of item $i$.
The probability $P(r|s)$ of observing a specific full ranking (permutation) $r=(r(1),r(2),\dots,r(K))$, where $r$ is a reordering of $\{1,\dots,K\}$, is defined sequentially based on Luce’s Choice Axiom [Luce1977TheYears]. Specifically, the probability is the product of the conditional probabilities of choosing item $r(i)$ from the set of items $R_{i}$ not yet ranked at step $i$:
@@
Figure 8: Global Procedural Progression. We visualize the temporal progression score averaged across all 80 Cholec80 [Twinanda2017EndoNet:Videos] videos, where higher scores indicate earlier phases. To generate this plot, we extract [CLS] embeddings using the frozen backbone of our trained PL-Stitch model and predict frame-wise scores using its temporal head. These scores are normalized per video to a unit scale ($0\to 1$) before computing the mean (solid line) and standard deviation (shaded region). Despite being trained on the different dataset, LEMON [Che2025LEMON:Settings], the model successfully generalizes to the unseen Cholec80 videos. The predicted score consistently decreases during the active surgical workflow. The observed deviations at the boundaries ($t<0.1$ and $t>0.9$) correspond to the camera entering and exiting the body, effectively marking the non-operative transitions surrounding the procedure.
This compels the encoder to assign a significantly higher score $s_{r^{*}(i)}$ to the correct frame (or patch) chosen at step $i$ compared to the log-sum-exponent of all remaining items. This probabilistic approach scales the error penalty proportionally to the ranking mistake severity, proving more robust than permutation classification [Misra2016ShuffleVerification, Wei2018LearningTime, Xu2019Self-SupervisedPrediction, Fernando2017Self-SupervisedNetworks] or pairwise losses [Hu2021ContrastLearning, Lee2017UnsupervisedSequences, Sermanet2017Time-ContrastiveObservation].
@@ A.3 Image Branch: Spatio-temporal Jigsaw
The Image Branch implements the $\mathcal{L}_{\text{\text{jigsaw}}}$ objective, another instantiation of the $\mathcal{L}_{\text{PL}}$ loss, designed to capture fine-grained object correspondence using local temporal context, inspired by  [Gupta2023SiameseAutoencoders, Liu2025WhenLearning].
@@ A.4 Learned Global Workflow Progression
To investigate whether PL-Stitch captures the global procedural structure, we conducted a qualitative analysis on the Cholec80 [Twinanda2017EndoNet:Videos] dataset. We employed the model pretrained on the large-scale LEMON dataset [Che2025LEMON:Settings], which contains no samples from Cholec80, ensuring a strict zero-shot evaluation. Specifically, we processed all 80 downstream videos by extracting [CLS] embeddings via the frozen PL-Stitch backbone and subsequently generating frame-wise real-valued scores using the pretrained temporal head.
@@ B.1 Pretraining Details
  • Backbone and Optimizer: We use a ViT-B/16 [Dosovitskiy2020AnScale] backbone. The optimizer is AdamW, with a base learning rate of $4\times 10^{-4}$, a weight decay of $0.05$, and a total batch size of 240.
  • Pretraining Length: Surgical models are pretrained on the LEMON dataset [Che2025LEMON:Settings] for 30 epochs. Cooking models are pretrained directly on their respective official training sets for 100 epochs. These durations were specifically selected to guarantee that the models are trained sufficiently well and that the training loss has effectively converged for each domain.
@@
  • $\mathcal{L}_{\text{MIM}}$ Parameters: A block-wise masking ratio of $30\%$ is applied to the current frame $v_{t}$, following the iBOT protocol [Zhou2022IBOT:Tokenizer].
  • $\mathcal{L}_{\text{jigsaw}}$ parameters: Following [Liu2025WhenLearning], which samples context frames from temporal offsets in the range $\pm[0.15T,0.25T]$ on Kinetics-400 [Carreira2017QuoDataset] videos (average duration $T$ $\approx 10$ s), we randomly sample the temporal offsets $\tau_{1}$ and $\tau_{2}$ from the range $[1.5,2.5]$ s relative to the current frame.
@@ C.1 Evaluation Datasets
  • Cholec80 [Twinanda2017EndoNet:Videos]: This dataset consists of 80 videos of laparoscopic cholecystectomy procedures. The objective is surgical phase recognition, a frame-wise classification task where the model must assign one of 7 distinct surgical phases (e.g., Preparation, Calot-Triangle Dissection, Clipping and Cutting) to every frame in the video.
  • AutoLaparo [Wang2022AutoLaparo:Hysterectomy]: A dataset containing 21 videos of laparoscopic hysterectomy procedures. The task is surgical phase recognition, challenging the model to identify the current phase from 7 defined surgical phases across long, untrimmed videos.
  • M2CAI16 [Stauder2016TheChallenge]: This dataset features 41 videos of laparoscopic cholecystectomy. The task is surgical phase recognition, where the model must recognize 8 surgical phases.
  • Breakfast [Kuehne2014TheActivities]: A dataset comprising 1712 videos capturing 10 different breakfast preparation activities (e.g., making coffee, pancakes) from a third-person perspective. The task is temporal action segmentation, which involves densely classifying video frames into 48 fine-grained action steps (e.g., “pour milk”, “crack egg”).
  • GTEA [Fathi2011LearningActivities]: An egocentric (first-person) dataset containing 28 videos of daily cooking activities. The task is temporal action segmentation across 11 action classes, presenting unique challenges due to severe camera motion and hand occlusions typical of wearable cameras.
@@ C.2 Evaluation Metrics
VideoMAEv2 [Wang2023VideoMAEMasking] $50.3\pm 1.9$ $56.5\pm 1.7$ $51.4\pm 3.4$
DINO [Caron2021EmergingTransformers] $75.5\pm 1.6$ $73.5\pm 1.5$ $68.9\pm 2.3$
iBOT [Zhou2022IBOT:Tokenizer] $75.6\pm 2.2$ $75.9\pm 1.2$ $71.5\pm 2.1$
@@
k-NN evaluation. Following established self-supervised learning benchmarks [Caron2021EmergingTransformers, Zhou2022IBOT:Tokenizer, Oquab2023DINOv2:Supervision], we assess the quality of the frozen feature space using non-parametric $k$-Nearest Neighbor ($k$-NN) classification. We first extract features from the frozen backbone for all frames in the training and testing sets. For each test frame, we retrieve its $k=20$ nearest neighbors from the training set based on cosine similarity. The predicted class is determined via weighted majority voting, where neighbors are weighted by their similarity scores. This protocol provides a direct measure of the semantic separability of the learned representation without requiring parameter updates. We report the top-1 accuracy.

### v2

@@ 3.5 Image Branch: Local and Spatio-temporal Learning
Masked Image Modeling. We adopt the masked image modeling (MIM) objective from iBOT [64] on the current frame $v_{t}$ to establish a robust frame-level feature representation, optimized with the loss $\mathcal{L}_{\text{MIM}}$.
@@ 4 Experiments
1 Pairwise [20] $75.8$ $75.4$
2 Permutation (CE) [31] $74.5$ $70.1$
@@ 4.1 Experimental Setup
Datasets and Evaluation Protocols. We evaluate on five widely used datasets from the surgical and cooking domains, selected for their long-range, multi-step procedural workflows with fixed sequential dependencies. For surgical procedures, we evaluate the temporal phase recognition task using AutoLaparo [54] (laparoscopic hysterectomy), along with Cholec80 [47] and M2CAI16 [41] (both laparoscopic cholecystectomy). For cooking activities, we evaluate the temporal action segmentation task using GTEA [12] (egocentric cooking) and Breakfast [24] (third-person cooking). We assess feature quality using two standard protocols: Linear Probing, which trains a linear classifier on frozen features, and k-NN Classification [3], which uses $k=20$ nearest neighbors in the feature space for non-parametric evaluation. For k-NN, we report top-1 accuracy. For Linear probing, we follow established benchmarks to report frame-wise accuracy and F1-score for surgical datasets [6, 54, 28], and frame-wise accuracy, segmental edit distance, and the segmental F1 score at overlapping thresholds 10%, 25% and 50% for cooking datasets [26, 12, 24].
@@
Pretraining Details. We adopt ViT-B/16 [11] as our framework’s backbone. For our temporal ranking objective ($\mathcal{L}_{\text{vid}}$), we sample k=8 frames from each video of length $T$. We first select a random starting frame $t_{0}$ and then a random, uniform step length $\Delta t$, such that $t_{0}+7\Delta t\leq T$. Videos are resampled multiple times per epoch, providing broad coverage of the procedural structure. For the $\mathcal{L}_{\text{jigsaw}}$ objective, the past and future context frames are sampled from the temporal ranges [-2.5, -1.5]s and [+1.5, +2.5]s relative to the current frame, following [29]. For the $\mathcal{L}_{\text{MIM}}$ objective, we apply block-wise masking with a ratio of 30% as in [64]. The loss weights (Eq. 5) are set to $\lambda_{1}=1$, $\lambda_{2}=1$, and $\lambda_{3}=0.4$ with analysis in supplementary material.
Optimization Strategies. We employ different pretraining optimization strategies based on the target domain. For surgical datasets, models are pretrained for 30 epochs on the large-scale surgical video dataset LEMON [6] to leverage its domain-specific knowledge. We use AdamW with a base learning rate of $4\times 10^{-4}$, a 3-epoch cosine warmup, and a total batch size of 240. For cooking datasets, we pretrain directly on their respective official training sets for 100 epochs with 10 warmup epochs, as no comparable large-scale pretraining dataset exists for this highly variable domain; other optimizer settings are kept identical. Experiments utilize 4 NVIDIA A100 (40GB) GPUs and PyTorch 2.5.1 [33].
Figure 7: Attention maps queried by the [CLS] token comparing our method, PL-Stitch, against DINO, iBOT, and VideoMAEv2 on a sequence of frames from AutoLaparo [54] during the Dividing Ligament and Peritoneum phase. The light blue box indicates the instrument operating area, which is critical for understanding the surgical phase. While other methods produce diffuse attention that changes erratically over time, our PL-Stitch maintains a consistent and accurate focus on the target area (the instrument and its operating area).
@@ 4.2 Main Quantitative Results
Surgical Phase Recognition. As shown in Table 1, PL-Stitch consistently outperforms all baselines on AutoLaparo, Cholec80, and M2CAI16 under both linear probing and k-NN evaluation. This demonstrates superior feature quality over both generalist (G) and specialist (S) methods. The performance gains are particularly striking in the k-NN evaluation, which directly assesses feature space quality. On Cholec80, PL-Stitch achieves 81.7% k-NN accuracy, a significant +11.4 pp gain over the strong iBOT baseline [64]. This trend holds for AutoLaparo (+7.2 pp) and M2CAI16 (+9.1 pp) as well. Furthermore, PL-Stitch also achieves the highest scores in all linear probing metrics, underscoring the discriminative power of its frozen features.
@@ 4.3 Ablation Study
We conduct ablations on the Cholec80 phase recognition task by pretraining on LEMON [6] and report top-1 accuracy under linear probing and k-NN (k = 20). We use a ViT-S/16 backbone to reduce computational cost. Our default settings in the Tables 34, and 5 are highlighted in gray.
Effect of PL-Stitch Components. We analyze the contribution of each objective in Table 3. Our baseline (row 1), with only the $\mathcal{L}_{\text{MIM}}$ objective [64], achieves 69.4% k-NN accuracy. Adding our core temporal ranking objective ($\mathcal{L}_{\text{vid}}$) (row 2) provides the largest single boost, improving k-NN accuracy by 9.5 pp to 78.9%. Adding only the spatio-temporal jigsaw objective ($\mathcal{L}_{\text{jigsaw}}$) (row 3) provides a smaller benefit. Finally, our full PL-Stitch model (row 4), combining all three objectives, achieves the highest accuracy at 80.2%. This demonstrates that our global temporal and local jigsaw objectives are complementary, and their joint optimization leads to the most robust representation.
Impact of PL Ranking Formulation. We compare our Plackett-Luce (PL) formulation against traditional ordering objectives (Table 4). We replace our $\mathcal{L}_{\text{vid}}$ with two common baselines: 1) a Pairwise Loss [20], which enforces the order of all pairs in the sequence, and 2) a Permutation Classification loss [31], which treats each shuffled permutation of the sequence as a distinct class. Both baselines perform worse than our PL objective, confirming that the probabilistic, listwise formulation of the PL model is key to learning a robust, progress-aware representation.
@@ 4.4 Qualitative Results
Feature Space Visualization. Fig. 6 visualizes the frozen Cholec80 features using t-SNE [48]. Baseline features from VideoMAEv2, DINO, and iBOT overlap heavily, showing no clear phase separation. In contrast, PL-Stitch forms distinct, well-separated clusters that align with ground-truth phases. Its clustering quality ($\text{ARI}:0.3536,\text{NMI}:0.4546$) is more than twice as high as the second-best baseline, DINO ($\text{ARI}:0.1043,\text{NMI}:0.2051$). This clear separation visually explains its high k-NN accuracy and confirms its learned procedural awareness.
@@ A.1 Plackett-Luce Ranking Formulation
Our framework leverages the Plackett-Luce (PL) model [35, 30] to structure both the global and local objectives as listwise ranking problems. The PL model is parameterized by a vector of positive real-valued scores $s=(s_{1},\dots,s_{K})$, where $K$ is the number of items being ranked. The score $s_{i}$ represents the overall utility or preference strength of item $i$.
The probability $P(r|s)$ of observing a specific full ranking (permutation) $r=(r(1),r(2),\dots,r(K))$, where $r$ is a reordering of $\{1,\dots,K\}$, is defined sequentially based on Luce’s Choice Axiom [30]. Specifically, the probability is the product of the conditional probabilities of choosing item $r(i)$ from the set of items $R_{i}$ not yet ranked at step $i$:
@@
Figure 8: Global Procedural Progression. We visualize the temporal progression score averaged across all 80 Cholec80 [47] videos, where higher scores indicate earlier phases. To generate this plot, we extract [CLS] embeddings using the frozen backbone of our trained PL-Stitch model and predict frame-wise scores using its temporal head. These scores are normalized per video to a unit scale ($0\to 1$) before computing the mean (solid line) and standard deviation (shaded region). Despite being trained on the different dataset, LEMON [6], the model successfully generalizes to the unseen Cholec80 videos. The predicted score consistently decreases during the active surgical workflow. The observed deviations at the boundaries ($t<0.1$ and $t>0.9$) correspond to the camera entering and exiting the body, effectively marking the non-operative transitions surrounding the procedure.
This compels the encoder to assign a significantly higher score $s_{r^{*}(i)}$ to the correct frame (or patch) chosen at step $i$ compared to the log-sum-exponent of all remaining items. This probabilistic approach scales the error penalty proportionally to the ranking mistake severity, proving more robust than permutation classification [31, 55, 58, 15] or pairwise losses [20, 25, 39].
@@ A.3 Image Branch: Spatio-temporal Jigsaw
The Image Branch implements the $\mathcal{L}_{\text{\text{jigsaw}}}$ objective, another instantiation of the $\mathcal{L}_{\text{PL}}$ loss, designed to capture fine-grained object correspondence using local temporal context, inspired by  [16, 29].
@@ A.4 Learned Global Workflow Progression
To investigate whether PL-Stitch captures the global procedural structure, we conducted a qualitative analysis on the Cholec80 [47] dataset. We employed the model pretrained on the large-scale LEMON dataset [6], which contains no samples from Cholec80, ensuring a strict zero-shot evaluation. Specifically, we processed all 80 downstream videos by extracting [CLS] embeddings via the frozen PL-Stitch backbone and subsequently generating frame-wise real-valued scores using the pretrained temporal head.
@@ B.1 Pretraining Details
  • Backbone and Optimizer: We use a ViT-B/16 [11] backbone. The optimizer is AdamW, with a base learning rate of $4\times 10^{-4}$, a weight decay of $0.05$, and a total batch size of 240.
  • Pretraining Length: Surgical models are pretrained on the LEMON dataset [6] for 30 epochs. Cooking models are pretrained directly on their respective official training sets for 100 epochs. These durations were specifically selected to guarantee that the models are trained sufficiently well and that the training loss has effectively converged for each domain.
@@
  • $\mathcal{L}_{\text{MIM}}$ Parameters: A block-wise masking ratio of $30\%$ is applied to the current frame $v_{t}$, following the iBOT protocol [64].
  • $\mathcal{L}_{\text{jigsaw}}$ parameters: Following [29], which samples context frames from temporal offsets in the range $\pm[0.15T,0.25T]$ on Kinetics-400 [4] videos (average duration $T$ $\approx 10$ s), we randomly sample the temporal offsets $\tau_{1}$ and $\tau_{2}$ from the range $[1.5,2.5]$ s relative to the current frame.
@@ C.1 Evaluation Datasets
  • Cholec80 [47]: This dataset consists of 80 videos of laparoscopic cholecystectomy procedures. The objective is surgical phase recognition, a frame-wise classification task where the model must assign one of 7 distinct surgical phases (e.g., Preparation, Calot-Triangle Dissection, Clipping and Cutting) to every frame in the video.
  • AutoLaparo [54]: A dataset containing 21 videos of laparoscopic hysterectomy procedures. The task is surgical phase recognition, challenging the model to identify the current phase from 7 defined surgical phases across long, untrimmed videos.
  • M2CAI16 [41]: This dataset features 41 videos of laparoscopic cholecystectomy. The task is surgical phase recognition, where the model must recognize 8 surgical phases.
  • Breakfast [24]: A dataset comprising 1712 videos capturing 10 different breakfast preparation activities (e.g., making coffee, pancakes) from a third-person perspective. The task is temporal action segmentation, which involves densely classifying video frames into 48 fine-grained action steps (e.g., “pour milk”, “crack egg”).
  • GTEA [12]: An egocentric (first-person) dataset containing 28 videos of daily cooking activities. The task is temporal action segmentation across 11 action classes, presenting unique challenges due to severe camera motion and hand occlusions typical of wearable cameras.
@@ C.2 Evaluation Metrics
VideoMAEv2 [50] $50.3\pm 1.9$ $56.5\pm 1.7$ $51.4\pm 3.4$
DINO [3] $75.5\pm 1.6$ $73.5\pm 1.5$ $68.9\pm 2.3$
iBOT [64] $75.6\pm 2.2$ $75.9\pm 1.2$ $71.5\pm 2.1$
@@
k-NN evaluation. Following established self-supervised learning benchmarks [3, 64, 32], we assess the quality of the frozen feature space using non-parametric $k$-Nearest Neighbor ($k$-NN) classification. We first extract features from the frozen backbone for all frames in the training and testing sets. For each test frame, we retrieve its $k=20$ nearest neighbors from the training set based on cosine similarity. The predicted class is determined via weighted majority voting, where neighbors are weighted by their similarity scores. This protocol provides a direct measure of the semantic separability of the learned representation without requiring parameter updates. We report the top-1 accuracy.